# Resolving the QCD Axion Domain Wall Problem with a Light Axion

### v1

@@ I Introduction
Various solutions have been proposed to resolve the domain wall problem of the QCD axion. One is to introduce an explicit PQ symmetry breaking term that generates a potential bias between different vacua, making the domain walls unstable and causing the network to collapse [17, 18]. However, this approach could spoil the PQ mechanism unless the CP phase of the explicit breaking is carefully chosen. Another approach is the Lazarides–Shafi (LS) mechanism [19, 20, 21], which identifies the degenerate vacua through gauge or global transformations, effectively reducing the domain wall number to unity. In this case, the axion strings decompose into $N_{\rm DW}$ (or more) hybrid or Alice strings, each attached to (at most) a single domain wall, thus solving the cosmological domain wall problem, because such a string-wall network collapses soon due to the wall tension. Nonetheless, realistic model-building along these lines, especially in the DFSZ framework, tends to require rather elaborate extensions.
@@
Recently, we investigated the dynamics of topological defects in the mixed axion system [44, 45], where various novel phenomena emerge compared to the single axion system. One is the formation of string bundles [36, 38], and the other is a transient bias, both of which play important roles in the dynamics of the axionic string-wall network. A string bundle refers to an isolated object in which strings associated with multiple axions are bound together by domain walls. It effectively corresponds to the cosmic string of a light axion at low energies. The transient bias arises when the potential of a light axion temporarily acts as a bias on the string-wall network before the onset of its oscillations.
@@ II Domain wall problem of the QCD axion
where $f_{a}$ is the decay constant of the QCD axion, and the radial degrees of freedom is omitted for simplicity. For our purpose it is sufficient to adopt this minimal model with a single PQ scalar. See Ref. [45] for the case of multiple PQ scalars.
@@
Once $m_{a}(T)$ or $m_{a0}$ becomes comparable to the Hubble parameter, the axion starts to roll down to the nearest potential minimum, and $N_{\rm DW}$ domain walls appear, attached to each string. If $N_{\rm DW}=1$, the network of the strings and walls collapses due to the wall tension and rapidly disappears. On the other hand, if $N_{\rm DW}>1$, the string-wall network is stable and approaches the so-called scaling regime, where each Hubble volume contains $\mathcal{O}(1)$ walls. In this regime, the energy density of domain walls decreases more slowly than matter and radiation, and eventually dominates the Universe. From the temperature anisotropies of the CMB, the domain wall tension is constrained as [10, 47]
@@
Since the domain wall tension is given by $\sigma_{a}=8m_{a}f_{a}^{2}/N_{\rm DW}^{2}$, which is larger than the QCD scale, this condition cannot be satisfied for the QCD axion. Thus, the post-inflationary scenario with $N_{\rm DW}>1$ suffers from the cosmological domain wall problem.
@@ III Solution with string bundles
where $f_{\phi}$ denotes the decay constant of $\phi$. In this section, we assume the post-inflationary scenario for both $a$ and $\phi$.
@@
When $S$ and $\Phi$ acquire nonzero vacuum expectation values (VEVs), two types of strings—$a$-strings and $\phi$-strings—are formed. Later, when the axion mass becomes comparable to the Hubble parameter, $m_{A}\sim H$, domain walls appear. These walls can be bounded by both types of strings. More specifically, each $a$-string is attached to $N_{\rm DW}$ domain walls, while each $\phi$-string is attached to a single wall.
Such a string-wall network subsequently transforms into string bundles consisting of a single $a$-string and $N_{\rm DW}$ $\phi$-anti-strings (or equivalently, a single $a$-anti-string and $N_{\rm DW}$ $\phi$-strings) [36, 45]. We show the sketch of string bundle formation in Fig. 1 (see also Fig. 5 in Appendix A). It is crucial that the domain wall number of $\phi$ is unity; if it were greater than one, the string-wall network would persist instead of collapsing into string bundles [45].
Figure 1: Schematic illustration of string bundle formation. The circles with $a$ and $\phi$ represent cosmic strings of $a$ and $\phi$, respectively, and the circles with $\bar{a}$ and $\bar{\phi}$ represent their corresponding anti-strings. The black lines connecting strings represent domain walls. Left panel: When the domain wall tension is too small to pull the strings together, the string-wall network stretches over cosmological scales. Right panel: Once the domain walls pull the strings together, strings and their anti-strings annihilate with each other, and a specific combination of strings forms a string bundle.
After the network has fully transitioned into string bundles, the domain walls are confined within the string bundles, and their energy density scales like that of cosmic strings. The resulting string bundle can be interpreted as a cosmic string associated with the massless axion $A_{L}$. As a result, only cosmic strings remain in the late Universe, without dominating the energy density. Interestingly, when the tension satisfies $\mu_{\phi}<(10^{14\text{--}15}\,\mathrm{GeV})^{2}$ [56, 57], current observations place no strong constraints. Thus, the domain wall problem is resolved. We will later discuss the preferred value of the tension that leads to the correct abundance of the QCD axion dark matter.
@@
Figure 2: $T_{\mathrm{DW}}$ as a function of $F$. The horizontal dashed line denotes $T_{\mathrm{DW}}=T_{\mathrm{QCD}}$. The behavior of $T_{\mathrm{DW}}$ changes on either side of this line.
@@
$\sigma_{A}(T)=8m_{A}(T)F^{2}\ .$ (15)
To form string bundles, the tension force of the domain walls must compete with that of the strings, so that they can pull the $\phi$-strings toward the $a$-strings. We expect that the domain wall tension force overcomes the string tension force when the energy density of domain walls exceeds that of strings. Assuming that both the string and wall networks are in the scaling regime, this condition becomes
@@
where $H$ is the Hubble parameter, and $\mu_{\phi}$ is the tension of $\phi$-strings. If this condition is satisfied at the time of domain wall formation, the string-wall network quickly transitions into string bundles. The typical size of a string bundle is given by $\mu_{\phi}/\sigma_{A}(T)$, which is of order the Hubble radius at formation. As the Universe cools, both $\sigma_{A}(T)$ and $1/H$ increase, causing the bundle size to become smaller than the Hubble radius at later times. The typical size of the string bundle is expected to be ${\cal O}(1-10)\mu_{\phi}/\sigma_{A}$, which is generally larger than the domain wall width $\sim 1/m_{A}$. We also obtain the same result by considering the repulsion force between the $\phi$-strings. [^1] Note that when the separation exceeds $1/m_{A}$, the strings interact only through the exchange of the massless axion. This analytical expectation is confirmed by our two-dimensional lattice simulations. See Appendix A for details.
On the other hand, if Eq. (16) is satisfied well after the domain wall formation, domain walls once settle into the scaling regime and then begin to collapse into string bundles when Eq. (16) is satisfied. In this case, we can estimate the temperature at the bundle formation by solving Eq. (16),
@@
[…]
@@
Let us estimate the QCD axion abundance in our scenario. If the string bundles are formed immediately after the domain wall formation (corresponding to the common diagonal line segment sloping downward from the upper left to the lower right in Fig. 3), the QCD axion is produced both via the misalignment mechanism and from partial annihilation of the string-wall network (see Fig. 1). Since the strings and walls do not completely annihilate, unlike in the standard scenario with $N_{\rm DW}=1$, the resulting axion abundance is expected to be slightly smaller than or comparable to the post-inflationary prediction based on a single PQ scalar model (see, e.g., Refs. [59, 60] for recent works).
On the other hand, if the string bundles are formed some time after the domain wall formation, QCD axions continue to be produced from strings and domain walls until the string bundle formation. After the formation of the string bundles, however, QCD axion production is expected to be kinematically suppressed. This is because the condition
@@
is likely satisfied in this case. One can see this by noting that $\sigma_{A}=8m_{A}F^{2}$, $\mu_{\phi}\sim\pi f_{\phi}^{2}$, and $F<f_{\phi}$ (see Appendix B for more details). Note that light axions continue to be emitted from the string bundles. Therefore, the dominant contribution to the QCD axion abundance arises from the string-wall network just before the string bundle formation.
Here, we follow the evaluation for the domain wall decay in the scaling regime [51]. In the scaling regime, the comoving area density of domain walls decreases, and their energy is emitted in the form of axion particles. We assume that the emitted axion has an energy of $\epsilon m_{A}$ with a constant parameter $\epsilon=\mathcal{O}(1)$. Since axion particles are diluted faster than domain walls in the scaling regime, we can neglect the contribution of axions emitted well before the bundle formation. Then, the axion number density after the bundle formation is given by
@@
where $R$ is the scale factor, and the subscript ‘bf’ denotes quantities at the bundle formation. Then, the energy density of the QCD axion in the later Universe is evaluated as
@@
=$=\begin{cases}\dfrac{3^{1/2}\times 5^{1/4}}{2^{1/4}\sqrt{\pi}}\dfrac{g_{*,\mathrm{bf}}^{3/4}}{\epsilon g_{*s,\mathrm{bf}}}\dfrac{\chi_{0}^{1/4}\sqrt{F\mu_{\phi}}}{M_{\mathrm{Pl}}^{3/2}}&(T_{\mathrm{bf}}<T_{\mathrm{QCD}})\\ \dfrac{1}{\epsilon g_{*s,\mathrm{bf}}}\left(\dfrac{2^{3n-2}\times 45^{n+2}g_{*,\mathrm{bf}}^{n+6}}{\pi^{2n+4}}\dfrac{\mu_{\phi}^{4}\chi_{0}^{n+2}F^{2n+4}}{M_{\mathrm{Pl}}^{2n+12}T_{\mathrm{QCD}}^{2n}}\right)^{\frac{1}{2n+8}}&(T_{\mathrm{bf}}\geq T_{\mathrm{QCD}})\end{cases}\ .$ (20)
[…]
@@
So far we have focused on the case of $N_{\rm DW}\neq 1$ as our primary motivation is to solve the domain wall problem of the QCD axion. However, it is worth noting that string bundles can also form in the case of $N_{\rm DW}=1$, provided that there is a light axion $\phi$ coupled to gluons.
@@
Let us now briefly comment on the LS mechanism [19]. In its canonical form, the model contains a gauged continuous symmetry. The color anomaly breaks the $U(1)_{\rm PQ}$ symmetry down to its subgroup $\mathbb{Z}_{N_{\rm DW}}$, which can be embedded in the center of the unbroken gauge group. As a result, the $N_{\rm DW}$ vacua are related by a gauge transformation and are therefore physically identical. That is to say, by introducing the gauge symmetry, the topology of vacua is modified such that there is no domain wall problem. In terms of topological defects, each axionic string is then attached to only a single domain wall, so the string–wall network is topologically unstable and annihilates quickly. In fact, a similar dynamics is possible without embedding the $\mathbb{Z}_{N_{\rm DW}}$ into center symmetry [20], in which case the axionic strings split into Alice strings. It is also possible to make use of a global symmetry, in which case the $N_{\rm DW}$ vacua are connected by a flat direction, i.e., a light or nearly massless NG mode [61, 62]. Our scenario has a similarity to this global-type LS mechanism in the sense that the $N_{\rm DW}$ vacua are connected by an additional light axion. Thus, the $N_{\rm DW}$ vacua correspond to the same physical vacuum plus a flat direction, so they can in principle be connected once an appropriate boundary of walls appears. Whether such a boundary forms depends on the domain-wall number carried by the $\phi$–string: if this number is unity the boundary is readily produced and the network reorganizes into stable string bundles; if it is two or larger the wall network forms before the boundary can develop, and the domain wall problem cannot be solved [45]. [^2] Note that the domain wall problem is mitigated in this case, since the tension can be much smaller than the QCD axion domain wall. Unlike the LS mechanism, where all topological defects disappear, our scenario leaves behind stable string bundles as the final configuration. This possibility was discussed in Ref. [63] in the context of the cosmic birefringence, but the detailed study in the context of resolving the QCD axion domain wall problem and the axion dark matter production has not yet been carried out.
@@
Our solution to the domain wall problem predicts the existence of a string-bundle network in the late-time Universe. This remnant string bundle serves as a smoking-gun signature of our scenario. If the massless axion couples to photons—as is often the case in UV completions333 This is indeed the case when $\phi$ couples to gluons as in the KSVZ axion model [54, 55]. —the stable string bundle induces anisotropic cosmic birefringence [64, 65]. Therefore, the simultaneous discovery of the QCD axion dark matter and anisotropic cosmic birefringence would provide strong evidence for this scenario. Alternatively, the light axion may be probed via future observations of gamma rays or X-rays through axion–photon conversion, particularly if the decay constant is smaller than $10^{10\text{--}11}\,\mathrm{GeV}$. Intriguingly, this mechanism might even explain certain observed events [66].

### v2

@@ I Introduction
Various solutions have been proposed to resolve the domain wall problem of the QCD axion. One is to introduce an explicit PQ symmetry breaking term that generates a potential bias between different vacua, making the domain walls unstable and causing the network to collapse [17, 18]. However, this approach could spoil the PQ mechanism unless the CP phase of the explicit breaking is carefully chosen. See also Refs. [19, 20, 21, 22] for cases where the PQ symmetry is dynamically broken. Another approach is the Lazarides–Shafi (LS) mechanism [23, 24, 25], which identifies the degenerate vacua through gauge or global transformations, effectively reducing the domain wall number to unity. In this case, the axion strings decompose into $N_{\rm DW}$ (or more) hybrid or Alice strings, each attached to (at most) a single domain wall, thus solving the cosmological domain wall problem, because such a string-wall network collapses soon due to the wall tension. Nonetheless, realistic model-building along these lines, especially in the DFSZ framework, tends to require rather elaborate extensions.
@@
Recently, we investigated the dynamics of topological defects in the mixed axion system [48, 49], where various novel phenomena emerge compared to the single axion system. One is the formation of string bundles [40, 42], and the other is a transient bias [49], both of which play important roles in the dynamics of the axionic string-wall network. A string bundle refers to an isolated object in which strings associated with multiple axions are bound together by domain walls. It effectively corresponds to the cosmic string of a light axion at low energies. The transient bias arises when the potential of a light axion temporarily acts as a bias on the string-wall network before the onset of its oscillations.
@@ II Domain wall problem of the QCD axion
where $f_{a}$ is the decay constant of the QCD axion, and the radial degrees of freedom are omitted for simplicity. For our purpose it is sufficient to adopt this minimal model with a single PQ scalar. See Ref. [49] for the case of multiple PQ scalars.
@@
Once $m_{a}(T)$ or $m_{a0}$ becomes comparable to the Hubble parameter, the axion starts to roll down to the nearest potential minimum, and $N_{\rm DW}$ domain walls appear, attached to each string. If $N_{\rm DW}=1$, the network of the strings and walls collapses due to the wall tension and rapidly disappears. On the other hand, if $N_{\rm DW}>1$, the string-wall network is stable and approaches the so-called scaling regime, where each Hubble volume contains $\mathcal{O}(1)$ walls. In this regime, the energy density of domain walls decreases more slowly than matter and radiation, and eventually dominates the Universe. From the observed temperature anisotropies of the CMB, the domain wall tension is constrained as [10, 51]
@@
Since the domain wall tension is given by $\sigma_{a}\simeq 8m_{a}f_{a}^{2}/N_{\rm DW}^{2}$, which is larger than the QCD scale, this condition cannot be satisfied for the QCD axion. Thus, the post-inflationary scenario with $N_{\rm DW}>1$ suffers from the cosmological domain wall problem [10].
@@ III Solution with string bundles
where $f_{\phi}$ denotes the decay constant of $\phi$, and the radial mode is omitted for simplicity. In this section, we assume the post-inflationary scenario for both $a$ and $\phi$.
@@
When $S$ and $\Phi$ acquire nonzero vacuum expectation values, two types of strings—$a$-strings and $\phi$-strings—are formed. Later, when the axion mass becomes comparable to the Hubble parameter, $m_{A}\sim H$, domain walls appear. These walls can be bounded by both types of strings. More specifically, each $a$-string is attached to $N_{\rm DW}$ domain walls, while each $\phi$-string is attached to a single wall.
Such a string-wall network subsequently transforms into string bundles consisting of a single $a$-string and $N_{\rm DW}$ $\phi$-anti-strings (or equivalently, a single $a$-anti-string and $N_{\rm DW}$ $\phi$-strings) [40, 42, 60, 49]. We show the sketch of string bundle formation in Fig. 1 (see also Fig. 6 in Appendix A). It is crucial that the domain wall number of $\phi$ is unity; if it were greater than one, the string-wall network would persist instead of collapsing into string bundles [49].
Figure 1: Schematic illustration of string bundle formation. The circles with $a$ and $\phi$ represent cosmic strings of $a$ and $\phi$, respectively, and the circles with $\bar{a}$ and $\bar{\phi}$ represent their corresponding anti-strings. The black lines connecting strings represent domain walls. Left panel: When the domain wall tension is too small to pull the strings together, the string-wall network stretches over cosmological scales. Right panel: Once the domain walls pull the strings together, the strings and their anti-strings annihilate with each other, and a specific combination of strings forms a string bundle.
After the network has fully transitioned into string bundles, the domain walls are confined within the string bundles, and their energy density scales like that of cosmic strings. The resulting string bundle can be interpreted as a cosmic string associated with the massless axion $A_{L}$. As a result, only cosmic strings remain in the late Universe, without dominating the energy density. Interestingly, when the tension satisfies $\mu_{\phi}<(10^{14\text{--}15}\,\mathrm{GeV})^{2}$ [61, 62, 63], current observations place no strong constraints. Thus, the domain wall problem is resolved. [^1] The same dynamics were pointed out in Ref. [64] in the context of connecting the cosmic birefringence to the QCD axion. We will later discuss the preferred value of the tension that leads to the correct abundance of the QCD axion dark matter.
@@
Figure 2: $T_{\mathrm{DW}}$ as a function of $F$. The horizontal dashed line denotes $T_{\mathrm{DW}}=T_{\mathrm{QCD}}$. The behavior of $T_{\mathrm{DW}}$ changes on either side of the dashed line.
@@
$\sigma_{A}(T)\simeq 8m_{A}(T)F^{2}\ .$ (15)
To form string bundles, the tension force of the domain walls must compete with that of the strings, so that they can pull the $\phi$-strings toward the $a$-strings. We expect that the domain wall tension force overcomes the string tension force when the energy density of domain walls exceeds that of strings. For simplicity, we focus on the case of $f_{a}/N_{\rm DW}\ll f_{\phi}$ so that the QCD axion mostly consists of $a$. In this case, the decay constant for the heavy axion (i.e., the QCD axion) is approximately given by $F\simeq f_{a}/N_{\rm DW}$. [^2] In the opposite case of $f_{a}/N_{\rm DW}\gg f_{\phi}$, the QCD axion mainly consists of $\phi$. For a given $f_{a}$, the string bundles start to form right after the domain walls appear, and the QCD axion abundance is considered to be suppressed compared to the case studied in the main text. Note that, in this case, there is no need to solve the domain wall problem from the beginning because $\phi$ has a unit domain wall number, but by adding an additional axion ($a$) one can similarly generate string bundles, which have the same interesting cosmological implications. Assuming that both the string and wall networks are in the scaling regime, this domain-wall dominance condition becomes
@@
where $H$ is the Hubble parameter, and $\mu_{\phi}(\sim\pi f_{\phi}^{2})$ is the tension of $\phi$-strings. We can estimate the temperature at which the bundles start to form by solving Eq. (16),
@@
We show the dependence of $T_{\mathrm{bf}}$ as a function of $F$ for $F<f_{\phi}/2$ with $\mu_{\phi}=\pi f_{\phi}^{2}$ in Fig. 3. In this region, $T_{\mathrm{bf}}$ is always smaller than $T_{\mathrm{DW}}$. For values of the decay constant in the range $10^{9}\,\mathrm{GeV}\lesssim F\lesssim 10^{12}\,\mathrm{GeV}$, where the QCD axion does not overclose the Universe and still satisfies astrophysical bounds, the string bundles are found to start forming at temperatures around $\mathcal{O}(0.01$$1)\,\mathrm{GeV}$ depending on $\mu_{\phi}$.
The typical size of a string bundle is given by ${\cal O}(1-10)\mu_{\phi}/\sigma_{A}(T)$, which is of order the Hubble radius at the bundle formation. As the Universe cools, both $\sigma_{A}(T)$ and $1/H$ increase, causing the bundle size to become smaller than the Hubble radius at later times. The typical size of the string bundle is generally larger than the domain wall width $\sim 1/m_{A}$. We also obtain the same result by considering the repulsion force between the $\phi$-strings. [^3] Note that when the separation exceeds $1/m_{A}$, the strings interact only through the exchange of the massless axion. This analytical expectation is confirmed by our two-dimensional lattice simulations. See Appendix A for details.
@@
Let us estimate the QCD axion abundance in our scenario. Since the string bundles are formed some time after the domain wall formation, QCD axions continue to be produced from strings and domain walls until the string bundle formation begins. After the formation of the string bundles begins, however, QCD axion production is expected to be kinematically suppressed. This is because the condition
@@
is satisfied in this case. One can see this by noting that $\sigma_{A}=8m_{A}F^{2}$, $\mu_{\phi}\sim\pi f_{\phi}^{2}$, and $F<f_{\phi}/2$ (see Appendix B for more details). Note that light axions continue to be emitted from the string bundles. Therefore, the dominant contribution to the QCD axion abundance arises from the string-wall network just before the string bundle formation begins.
Here, we follow the evaluation for the domain wall decay in the scaling regime [55]. In the scaling regime, the comoving area density of domain walls decreases, and their energy is emitted in the form of axion particles. We assume that the emitted axion has an energy of $\epsilon m_{A}$ with a constant parameter $\epsilon=\mathcal{O}(1)$. Since axion particles are diluted faster than domain walls in the scaling regime, we can neglect the contribution of axions emitted well before the beginning of the bundle formation. Then, the axion number density after the bundle formation is given by
@@
where $R$ is the scale factor, and the subscript ‘bf’ denotes quantities at $T=T_{\rm bf}$. Then, the energy density of the QCD axion in the later Universe is evaluated as
@@
$=\begin{cases}\dfrac{3^{1/2}\times 5^{1/4}}{2^{1/4}\sqrt{\pi}}\dfrac{g_{*,\mathrm{bf}}^{3/4}}{\epsilon g_{*s,\mathrm{bf}}}\dfrac{\chi_{0}^{1/4}\sqrt{F\mu_{\phi}}}{M_{\mathrm{Pl}}^{3/2}}&(T_{\mathrm{bf}}<T_{\mathrm{QCD}})\\ \dfrac{1}{\epsilon g_{*s,\mathrm{bf}}}\left(\dfrac{2^{3n-2}\times 45^{n+2}g_{*,\mathrm{bf}}^{n+6}}{\pi^{2n+4}}\dfrac{\mu_{\phi}^{4}\chi_{0}^{n+2}F^{2n+4}}{M_{\mathrm{Pl}}^{2n+12}T_{\mathrm{QCD}}^{2n}}\right)^{\frac{1}{2n+8}}&(T_{\mathrm{bf}}\geq T_{\mathrm{QCD}})\end{cases}\ .$ (20)
We show the abundance of the QCD axion produced by (partial) domain wall annihilation in Fig. 4 for $F=10^{9},\,10^{10},\,10^{11},\,10^{12}$ GeV. Here we set $\mu_{\phi}=\pi f_{\phi}^{2}$. One can see that, considering that the misalignment contribution becomes dominant for $F\gtrsim{\cal O}(10^{11})$ GeV, the preferred value of $f_{\phi}$ lies around $10^{13\text{--}15}$ GeV. [^4] The PQ symmetry of the scalar $\Phi$ can be restored even with a large decay constant if the reheating temperature is sufficiently high; it can also be restored at a lower reheating temperature or inflation scale when the $\Phi$ potential is flat, its nonminimal coupling to gravity is large, or $\Phi$ serves as the waterfall field in hybrid inflation. This suggests that such string bundles may be probed by future CMB observations, pulsar timing arrays such as SKA, and gravitational experiments such as DECIGO, BBO, and LISA (see Ref. [62]). Note that, if the string tension grows logarithmically, the associated CMB and gravitational-wave signals are enhanced. At large scales, string bundles can be approximated as ordinary global strings with an effective tension $N_{\rm DW}^{2}\mu_{\phi}$. [^5] This will be replaced by $\mu_{a}\sim\pi f_{a}^{2}$ if we impose the inverse hierarchy, $f_{a}\gg N_{\rm DW}f_{\phi}$, where $\phi$ mainly plays a role of the QCD axion. Our model therefore provides a concrete example of a global string that can act as a source of gravitational waves without accompanying domain walls. Massless axions $A_{L}$ are produced during the evolution of string bundles, but their energy density never dominates the universe; thus, there is no relevant constraint from $N_{\rm eff}$. On the other hand, the internal structure of the string bundles could make characteristic features on the gravitational wave signatures from the ordinary global string case. For the case with $f_{\phi}=10^{15}\,{\rm GeV}$, the typical width of the string bundle is $\sim 10^{6-7}\,{\rm m}$, when the QCD axion explains whole dark matter. The relative motion of $\phi$-strings could produce gravitational waves with the frequency $\sim 10^{1-2}\,{\rm Hz}$. More detailed study is left for future work.
On the other hand, for $F<10^{9}$ GeV, already subject to severe astrophysical bounds, the string bundle formation must be delayed in order to yield the right amount of the QCD axion dark matter. This requires a much larger $f_{\phi}$, which in turn implies a large string tension $\mu_{\phi}$ that is likely already excluded by current observational constraints. This issue can be circumvented if the string bundles subsequently annihilate, which can be achieved by introducing an additional potential as discussed below.
Figure 4: The QCD axion abundance from the string-wall network as a function of the decay constant $f_{\phi}$, for the QCD axion decay constant $F=10^{12},\,10^{11},\,10^{10}\,10^{9}$ GeV from top to bottom. The gray shaded region indicates where the abundance exceeds the observed dark matter abundance $\Omega_{c}h^{2}\simeq 0.12$.
@@
So far, we have focused on the case of $N_{\rm DW}\neq 1$ as our primary motivation is to solve the domain wall problem of the QCD axion. However, it is worth noting that string bundles can also form in the case of $N_{\rm DW}=1$, provided that there is a light axion $\phi$ coupled to gluons.
@@
Let us now briefly comment on the LS mechanism [23]. In its canonical form, the model contains a gauged continuous symmetry. The color anomaly breaks the $U(1)_{\rm PQ}$ symmetry down to its subgroup $\mathbb{Z}_{N_{\rm DW}}$, which can be embedded in the center of the unbroken gauge group. As a result, the $N_{\rm DW}$ vacua are related by a gauge transformation and are therefore physically identical. That is to say, by introducing the gauge symmetry, the topology of vacua is modified such that there is no domain wall problem. [^6] See also Ref. [66], which points out that even in models where the $\mathbb{Z}_{N_{\rm DW}}$ subgroup of $U(1)_{\rm PQ}$ is embedded in a continuous gauge group, the cosmological dynamics of string formation may still fail to eliminate domain walls, depending on the UV completion. In terms of topological defects, each axionic string is then attached to only a single domain wall, so the string–wall network is topologically unstable and annihilates quickly. In fact, a similar dynamics is possible without embedding the $\mathbb{Z}_{N_{\rm DW}}$ into center symmetry [24], in which case the axionic strings split into Alice strings. It is also possible to make use of a global symmetry, in which case the $N_{\rm DW}$ vacua are connected by a flat direction, i.e., a light or nearly massless NG mode [67, 68]. Our scenario has a similarity to this global-type LS mechanism in the sense that the $N_{\rm DW}$ vacua are connected by an additional light axion. Thus, the $N_{\rm DW}$ vacua correspond to the same physical vacuum plus a flat direction, so they can, in principle, be connected once an appropriate boundary of walls appears. Whether such a boundary forms depends on the domain-wall number carried by the $\phi$–string: if this number is unity, the boundary is readily produced and the network reorganizes into stable string bundles; if it is two or larger, the wall network forms before the boundary can develop, and the domain wall problem cannot be solved [49]. [^7] Note that the domain wall problem is mitigated in this case, since the tension can be much smaller than the QCD axion domain wall. Unlike the LS mechanism, where all topological defects disappear, our scenario leaves behind stable string bundles as the final configuration. This possibility was discussed in Ref. [64] in the context of the cosmic birefringence, but the detailed study in the context of resolving the QCD axion domain wall problem and the axion dark matter production has not yet been carried out.
@@
Our solution to the domain wall problem predicts the existence of a string-bundle network in the late-time Universe. This remnant string bundle serves as a smoking-gun signature of our scenario. If the massless axion couples to photons—as is often the case in UV completions [^8] This is indeed the case when $\phi$ couples to gluons as in the KSVZ axion model [58, 59]. —the stable string bundle induces anisotropic cosmic birefringence [69, 70]. Therefore, the simultaneous discovery of the QCD axion dark matter and anisotropic cosmic birefringence, as well as stochastic gravitational-wave backgrounds across multiple frequency bands-from the nanohertz range accessible to pulsar timing arrays to the millihertz–hertz range probed by space- and ground-based interferometers-would provide compelling evidence for this scenario. Alternatively, the light axion may be probed via future observations of gamma rays or X-rays through axion–photon conversion, particularly if the decay constant is smaller than $10^{10\text{--}11}\,\mathrm{GeV}$. Intriguingly, this mechanism might even explain certain observed events [71].